Explain why the spectroscopic term symbol for Boron in the ground state is 2P3/2.
1
Expert's answer
2020-05-27T10:32:32-0400

Electron configuration of Boron is "Z=1s^22s^22p^1"

For ground state"2p^1, s=\frac{1}{2} and" "l=1"

The possible value of J is "J=l+s=1+\frac{1}{2}=\frac{3}{2}"

And "L=l=1 ,L=1" this means "L=P"

Spectroscopic term symbol formula is

"^{2s+1}L_J=^{(2(\frac{1}{2})+1)}L_\frac{3}{2}=^2P_\frac{3}{2}"